owlshift 1-3-00


Subject: owlshift 1-3-00
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:28:48 CST


Split shift:

Earl Swallow 00:00 - 03:30
Hogan 3:30 - 07:00

No beam, but expect beam later today. MCR is trying to bring the complex
back on. Problems with the "Mubend QPM", NM2EU will not run up above
4kamps, and various other problems.

Quiet shift. Ran smoothly with cosmics for most of the shift. Had to do
one drt_fresh_start to clear a DDD3 veto.

Sent email to Dave Northacker about alcohol refill.

6 am -- RSK and EB arrives to work on BA

         EB installs new CsI constants based on run 15322 (first update of
         the new millenium )

ET is off. Should be back on this morning.

    -Hogan



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:28:50 CST